# Extremely Amenable Automorphism Groups of Countable Structures

### v1

@@ Abstract.
In this paper we address the question: How many pairwise non-isomorphic extremely amenable groups are there which are seprable metrizable or even Polish? We show that there are continuum many such groups. In fact we construct continuum many pairwise non-isomorphic extremely amenable groups as automorphism groups of countable structures. We also consider this classicfication problem from the point of view of descriptive set theory by showing that the class of all extremely amenable closed subgroups of $S_{\infty}$ is Borel and by obtaining some lower bounds for their isomorphism relation in the Borel reducibility hierarchy.
The second and the third authors acknowledge the partial support of their research by the National Natural Science Foundation of China (NSFC) grants 12250710128 and 12271263.
@@ 1. Introduction
Ever since Mitchell [17] asked in 1970 about the existence of non-trivial extremely amenable topological groups, many examples of such groups have been discovered.
@@
A third class of topological groups known to be extremely amenable in the literature comes from automorphism groups of countable structures. Pestov [24] gave the first example of such a group, which is ${\rm Aut}(\mathbb{Q},<)$, the automorphism group of a countable dense linear order without endpoints. Motivated by the idea in [24], Kechris–Pestov–Todorcevic [15] gave a complete characterization of when the automorphism group of a countable structure is extemely amenable, which is now called the KPT correspondence. The KPT correspondence connects the extreme amenability to the so-called structural Ramsey property of Fraïssé classes and gives a number of examples of extremely amenable groups. Through the work of Nešetřril–Rödl [19] [20] and Nešetřil [18] on Ramsey classes, we now have extremely amenable groups of the form $\mbox{\rm Aut}(M)$, where $M$ is the Fraïssé limit of a Fraïssé class of countable structures. Such structures include ordered graphs, ordered $K_{n}$-free graphs, ordered $\mathbb{Q}$-metric spaces, and ordered $\mathbb{N}$-metric spaces, etc.
In this paper we are concerned with the question: How many pairwise non-isomorphic extremely amenable groups are there? The only result that explicitly addresses this question is a theorem of Pestov in [24], which states that for each infinite cardinal $\kappa$, there is an extremely amenable topological group of weight $\kappa$. Since the weight is an isomorphism invariant, we obtain many pairwise non-isomorphic extremely amenable groups. However, we would like to study this question in the context of more familiar groups, in particular separable metrizable groups and Polish groups. The theorem of Pestov gives only one such example. Despite the diversity of examples mentioned above, many are seprable metrizable and even Polish, we found no results in the literature that differentiate the isomorphism types of the examples, and therefore some work needs to be done before one can claim that there are infinitely many pairwise non-isomorphic extremely amenable groups that are separable metrizable or Polish.
@@
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable iff $\inf\Delta=0$.
@@
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is Polish iff $\inf\Delta>0$.
@@
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Iso}(\mathbb{U}_{\Lambda})$ are isomorphic as topogoical groups.
@@
We then study the case of Polish groups. Here we consider the Fraïssé class of all finite ordered $\Delta$-metric spaces and its Fraïssé limit $\mathbb{U}^{<}_{\Delta}$. Regarding $\mathbb{U}^{<}_{\Delta}$ as a countable structure, the automorphism group $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$, equipped with the topology of pointwise convergence where the underlying space has the discrete topology, is a Polish group. In fact, it is isomorphic to a closed subgroup of $S_{\infty}$, the infinite permutation group. We use the following consequence of a theorem of Nešetřil [18] (for a more explicit restatement of Nešetřil’s theorem, see Theorem 13 of Nguyen Van Thé [22]).
@@
Let $\Delta$ be a countable distance value set. Then $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ is extremely amenable.
@@
    $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ and $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Lambda})$ are isomorphic as topological groups.
@@
There are continnum many pairwise non-isomorphic extremely amenable Polish groups $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$.
@@
In addition, we consider some related questions from the point of view of descriptive set theory. Following a program started by Kechris–Nies–Tent [14], we show that the class of extremely amenable closed subgroups of $S_{\infty}$ is a Borel class, and that the isomorphism relation on this class is an equivalence relation that is not smooth. Using some recent work of Calderoni–Marker–Motto Ros–Shani [2], we prove that a non-optimal lower bound for this equivalence relation in the Borel reducibility hierarchy is $=^{+}$, the equality relation of countable sets of real numbers. From the point of view of descriptive set theory, this is a strengthening of Corollary 1.6.
@@
The equality relation of countable sets of real numbers, $=^{+}$, is strictly Borel reducible to the isomorphism relation among extremely amenable Polish groups that are isomorphic to closed subgroups of $S_{\infty}$.
The rest of the paper is organized as follows. In Section 2 we define the basic concepts and notions and give a review of the results we need to cite in this paper. In Section 3 we prove Theorems 1.1 and 1.4. In Section 4 we study the notion of equivalence between distance value sets, and prove the first part of Theorem 1.3. In Section 5 we prove Theorems 1.2 and 1.5, which will allow us to derive the second part of Theorem 1.3 and Corollary 1.6. In Section 6 we prove that the class of extremely amenable closed subgroups of $S_{\infty}$ is Borel and state some consequences of our results about the isomorphism relation of extremely amenable groups. Finally, in Section 7 we make some concluding remarks and state some open questions.
Acknowledgment. We thank Filippo Calderoni, Ruiwen Li, Marcin Sabok, Assaf Shani, and Simon Thomas for their inspiring comments on an earlier version of the paper.
@@ 2. Preliminaries
In this section we review the basic concepts and notation used in the rest of the paper.
@@ 2.1. Topological dynamics
Our topological spaces and topological groups are all Hausdorff. Let $G$ be a topological group and let $X$ be a compact Hausdorff space. Let $a\colon G\times X\to X$ be a continuous action of $G$ on $X$. For $g\in G$ and $x\in X$, we also write $g\cdot x$ for $a(g,x)$. We say that $X$ is a $G$-flow. For $x\in X$, the orbit of $x$, denoted $[x]_{G}$, is the set $\{g\cdot x\colon g\in G\}$. We say that the $G$-flow $X$ is minimal if for every $x\in X$, $[x]_{G}$ is dense in $X$.
@@
A topological group $G$ is extremely amenable if every $G$-flow $X$ has a fixed point, i.e., there is $x\in X$ such that $g\cdot x=x$ for all $g\in G$. Equivalently, a topological group $G$ is extremely amenable iff the universal minimal $G$-flow is a singleton. By a theorem of Veech [31], no non-trivial locally compact topological groups can be extremely amenable. Thus extreme amenability is a largeness property for topological groups.
@@ 2.2. Fraïssé Theory
    a collection $(F_{j})_{j\in J}$ of function symbols; for each $j\in J$, there is an associate arity $m(j)$, which is a natural number, such that $F_{j}$ is an $m(j)$-ary function symbol.
The 00-ary function symbols, if any, are also called constant symbols. We say $\mathcal{L}$ is countable if $I$ and $J$ are countable.
@@
    $A$ is a set known as the universe of $\mathcal{A}$ and denoted as $|\mathcal{A}|$;
@@
If there is an embedding from $\mathcal{A}$ to $\mathcal{B}$, we say $\mathcal{A}$ is embeddable to $\mathcal{B}$, denoted by $\mathcal{A}\leq\mathcal{B}$. Finally, if $f$ is bijective, we say it is an isomorphism and if there is an isomorphism between $\mathcal{A}$ and $\mathcal{B}$, we say they are isomorphic, denoted by $\mathcal{A}\cong\mathcal{B}$.
@@ 2.3. The KPT correspondence
A Polish space is a separable, completely metrizable topological space. A Polish group is a topological group whose topology is Polish. An example of a Polish group is the infinite permutation group $S_{\infty}$. $S_{\infty}$ consists of all permutations of $\mathbb{N}$, where the group operation is composition. The topology on $S_{\infty}$ is the pointwise convergence topology, where the topology on $\mathbb{N}$ is discrete. Equivalently, $S_{\infty}$ has the subspace topology inherited from $\mathbb{N}^{\mathbb{N}}$, where $\mathbb{N}^{\mathbb{N}}$ has the product topology. Both the topologies on $\mathbb{N}^{\mathbb{N}}$ and on $S_{\infty}$ are Polish.
@@
A Fraïssé class $\mathcal{K}$ is said to be ordered if there is a binary relation symbol $<$ in $\mathcal{L}$ and, in every structure in $\mathcal{K}$, the interpretaton of $<$ is a linear order. The KPT correspondence unravels a deep connection between the property of a Fraïssé ordered class $\mathcal{K}$ and the extreme amenability of $\mbox{\rm Aut}(\mbox{Flim}\,\mathcal{K})$. This property is called the (structural) Ramsey property, which we define below.

### v2

@@ Abstract.
In this paper we address the question: How many pairwise non-isomorphic extremely amenable groups are there which are separable metrizable or even Polish? We show that there are continuum many such groups. In fact we construct continuum many pairwise non-isomorphic extremely amenable groups as automorphism groups of countable structures. We also consider this classification problem from the point of view of descriptive set theory by showing that the class of all extremely amenable closed subgroups of $S_{\infty}$ is Borel and their isomorphism relation is more complex than any isomorphism relation of countable structures in the Borel reducibility hierarchy.
2020 Mathematics Subject Classification:
22F05, 54H05
S. Gao and the F. Li acknowledge the partial support of their research by the National Natural Science Foundation of China (NSFC) grants 12271263 and 12250710128. R. Li acknowledges the partial support of his research by the NSFC grant 124B2001. S. Gao, F. Li and R. Li attended the Tianyuan Workshop on Computability Theory and Descriptive Set Theory in June 2025 when a part of this research was conducted, and we acknowledge the support by the Tianyuan Mathematics Research Center.
@@ 1. Introduction
A topological group $G$ is extremely amenable if for every continuous action of $G$ on a compact Hausdorff space $X$, there is a fixed point $x\in X$, i.e., for all $g\in G$, $g\cdot x=x$. This is a remarkable property for a topological group to have. In fact, Veech [35] proved that no locally compact topological groups can be extremely amenable. This explained why extremely amenable groups were not present in the early development of topological group theory, for instance in the study of Hilbert’s fifth problem.
Ever since Mitchell [20] asked in 1970 about the existence of non-trivial extremely amenable topological groups, many examples of such groups have been discovered. At first, these groups were thought to be pathological since they had to be obtained via intricate constructions. However, as time goes by, more familiar groups have been shown to be extremely amenable. The extremely amenable groups we now know of can be loosely sorted into three classes as follows.
@@
A third class of topological groups known to be extremely amenable in the literature comes from automorphism groups of countable structures. Pestov [28] gave the first example of such a group, which is ${\rm Aut}(\mathbb{Q},<)$, the automorphism group of a countable dense linear order without endpoints. Motivated by the idea in [28], Kechris–Pestov–Todorcevic [17] gave a complete characterization of when the automorphism group of a countable structure is extemely amenable, which is now called the KPT correspondence. The KPT correspondence connects the extreme amenability to the so-called structural Ramsey property of Fraïssé classes and gives a number of examples of extremely amenable groups. Through the work of Nešetřil–Rödl [22] [23] and Nešetřil [21] on Ramsey classes, we now have extremely amenable groups of the form $\mbox{\rm Aut}(M)$, where $M$ is the Fraïssé limit of a Fraïssé class of countable structures. Such structures include ordered graphs, ordered $K_{n}$-free graphs, ordered $\mathbb{Q}$-metric spaces, and ordered $\mathbb{N}$-metric spaces, etc. The KPT correspondence has also been generalized to the context of continuous logic by Melleray–Tsankov [19].
Despite the diversity of these examples, not much was known about the number of distinct isomorphism types of extremely amenable groups. Among the above examples, only finitely many pairwise non-isomorphic extremely amenable groups can be identified. This motivates us to study the following question: How many pairwise non-isomorphic extremely amenable groups are there?
The only result that explicitly addressed this question in the literature is a theorem of Pestov in [28], which states that for each infinite cardinal $\kappa$, there is an extremely amenable topological group of weight $\kappa$. Since the weight is an isomorphism invariant, we obtain many pairwise non-isomorphic extremely amenable groups. However, we would like to study this question in the context of more familiar groups, in particular separable metrizable groups and Polish groups. The theorem of Pestov gives only one such example.
Our main results of the paper confirm that there are continuum many non-isomorphic extremely amenable groups which are either Polish or separable metrizable. Our study focuses on the third class of examples in the above review.
@@
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable if and only if $\inf\Delta=0$.
@@
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is Polish if and only if $\inf\Delta>0$.
@@
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Iso}(\mathbb{U}_{\Lambda})$ are isomorphic as topological groups.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Our study of the isomorphism types of these extremely amenable groups goes much beyond counting the number of them. Another major effort we put in this paper is to study the isomorphism relation among these extremely amenable groups from the point of view of descriptive set theory. (For an overview of the descriptive set theory of equivalence relations, see Section 2). In view of Theorems 1.2 and 1.5, this is done by studying the complexity of the equivalence relation between countable distance value sets in the Borel reducibility hierarchy. We obtain the following penetrating results.
@@
  1. (1)
    The equivalence between countable distance value sets $\Delta$ with $\inf\Delta>0$ and $\sup\Delta<\infty$ is Borel bireducible with the universal $S_{\infty}$-orbit equivalence relation.
  2. (2)
    The equivalence between countable distance value sets $\Delta$ with $\inf\Delta>0$ and $\sup\Delta=\infty$ is Borel bireducible with the equivalence relation $^{+}$.
  3. (3)
    The equivalence between countable distance value sets $\Delta$ with $\inf\Delta=0$ is strictly in between $^{+}$ and $^{++}$ in the Borel reducibility hierarchy.
These results significantly strengthen Theorem 1.3 and Corollary 1.6. Theorem 1.7 (3) is motivated by and related to the recent work of Calderoni–Marker–Motto Ros–Shani [2].
Finally, inspired by a research program initiated by Kechris–Nies–Tent [16], we show that the class of extremely amenable closed subgroups of $S_{\infty}$ is a Borel class (this was also independently proved by Iwanow–Majcher-Iwanow (unpublished); see [24, Section 3]), which implies that the isomorphism relation among these groups is an analytic equivalence relation. Regarding its complexity in the Borel reducibility hierarchy, our results yield the following.
Theorem 1.8.
The isomorphism relation among extremely amenable Polish groups that are isomorphic to closed subgroups of $S_{\infty}$ is more complex than the universal $S_{\infty}$-orbit equivalence relation. In particular, it is analytic and non-Borel.
The rest of the paper is organized as follows. In Section 2 we define the basic concepts and notions and give a review of the results we need to cite in this paper. In particular, we review the descriptive set theory of equivalence relations. In Section 3 we prove Theorems 1.1 and 1.4. In Section 4 we prove Theorems 1.2 and 1.5. In Section 5 we prove Theorem 1.7. Finally, in Section 6 we prove that the class of extremely amenable closed subgroups of $S_{\infty}$ is Borel, derive Theorem 1.8, and state some questions left open by this paper.
Acknowledgment. We thank Filippo Calderoni, Wei Dai, Aleksander Iwanow, Andre Nies, Marcin Sabok, Assaf Shani, Simon Thomas, and Víctor Hugo Yañez for their inspiring comments on an earlier version of the paper.
@@ 2. Preliminaries
In this section we review some of the basic concepts and notation we will use in the rest of the paper.
@@ 2.1. Topological dynamics
Our topological spaces and topological groups are all Hausdorff. Let $G$ be a topological group and let $X$ be a compact Hausdorff space. Let $a\colon G\times X\to X$ be a continuous action of $G$ on $X$. For $g\in G$ and $x\in X$, we also write $g\cdot x$ for $a(g,x)$. We say that $X$ is a $G$-flow. For $x\in X$, the orbit of $x$, denoted $x]_{G}$, is the set $\{g\cdot x\colon g\in G\}$. We say that the $G$-flow $X$ is minimal if for every $x\in X$, $x]_{G}$ is dense in $X$.
@@
A topological group $G$ is extremely amenable if every $G$-flow $X$ has a fixed point, i.e., there is $x\in X$ such that $g\cdot x=x$ for all $g\in G$. Equivalently, a topological group $G$ is extremely amenable if and only if the universal minimal $G$-flow is a singleton. By a theorem of Veech [35], no non-trivial locally compact topological group can be extremely amenable. Thus extreme amenability is a largeness property for topological groups.
@@ 2.2. Fraïssé Theory
    a collection $(F_{j})_{j\in J}$ of function symbols; for each $j\in J$, there is an associated arity $m(j)$, which is a natural number, such that $F_{j}$ is an $m(j)$-ary function symbol.
The 0-ary function symbols, if any, are also called constant symbols. We say $\mathcal{L}$ is countable if $I$ and $J$ are countable.
@@
    AA is a set known as the universe of $\mathcal{A}$ and denoted as $|\mathcal{A}|$;
@@
If there is an embedding from $\mathcal{A}$ to $\mathcal{B}$, we say $\mathcal{A}$ is embeddable to $\mathcal{B}$, denoted by $\mathcal{A}\leq\mathcal{B}$. Finally, if $f$ is bijective, we say it is an isomorphism; if there is an isomorphism between $\mathcal{A}$ and $\mathcal{B}$, we say they are isomorphic, and denote $\mathcal{A}\cong\mathcal{B}$.
@@ 2.3. The KPT correspondence
A Polish space is a separable, completely metrizable topological space. A Polish group is a topological group whose topology is Polish. An example of a Polish group is the infinite permutation group $_{\infty}. SS_{\infty}$ consists of all permutations of $\mathbb{N}$, where the group operation is composition. The topology on $S_{\infty}$ is the pointwise convergence topology, where the topology on $\mathbb{N}$ is discrete. Equivalently, $S_{\infty}$ has the subspace topology inherited from $\mathbb{N}^{\mathbb{N}}$, where $\mathbb{N}^{\mathbb{N}}$ has the product topology. Both the topologies on $\mathbb{N}^{\mathbb{N}}$ and on $S_{\infty}$ are Polish.
@@
A Fraïssé class $\mathcal{K}$ is said to be ordered if there is a binary relation symbol << in $\mathcal{L}$ and, in every structure in $\mathcal{K}$, the interpretaton of << is a linear order. The KPT correspondence unravels a deep connection between the property of a Fraïssé ordered class $\mathcal{K}$ and the extreme amenability of $\mbox{\rm Aut}(\mbox{Flim}\,\mathcal{K})$. This property is called the (structural) Ramsey property, which we define below.
@@
2.4. Descriptive set theory of equivalence relations
In this paper we will be discussing the complexity of equivalence relations in the framework of descriptive set theory. We review the basic concepts and notions in this subsection. For other undefined terms and concepts, see [8].
Let $X$ be a set and let $\mathcal{B}$ be a $\sigma$-algebra of subsets of $X$. $(X,\mathcal{B})$ is a standard Borel space if there is a Polish topology $\tau$ on $X$ such that $\mathcal{B}$ is the $\sigma$-algebra of Borel sets generated (via iterations of countable union, countable intersection, and complementation) by $\tau$. Any Polish space is a standard Borel space. If $X$ is a standard Borel space and $Y$ is a Borel subset of $X$, then $Y$ is a standard Borel space. A subset $A$ of a standard Borel space is analytic if there is a standard Borel space $Y$ and there is a Borel function $f\colon Y\to X$ with $A=f(Y)$. An equivalence relation $E$ on a standard Borel space $X$ is Borel (or analytic) if $E$ is a Borel (analytic) subset of $X^{2}$.
For equivalence relations $E,F$ on standard Borel spaces $X,Y$, respectively, we say that $E$ is Borel reducible to $F$, denoted $E\leq_{B}F$, if there is a Borel function $f\colon X\to Y$ such that for all $x,x^{\prime}\in X$, we have
$xEx^{\prime}\iff f(x)Ff(x^{\prime}).$
We say that $E$ is Borel bireducible with $F$, denoted $E\sim_{B}F$, if both $E\leq_{B}F$ and $F\leq_{B}E$ hold. We say that $E$ is strictly Borel reducible to $F$, denoted $E<_{B}F$, if $E\leq_{B}F$ but $F\not\leq_{B}E$. The Borel reducibility preorder $\leq_{B}$ gives rise to a relative complexity hierarchy for equivalence relations on standard Borel spaces. Specifically, given an equivalence relation $E$, we attempt to determine its exact complexity by comparing it, in terms of Borel reducibility, with some benchmark equivalence relations previously studied and understood in the theory. If there is a benchmark equivalence relation $F$ so that we can show $E\sim_{B}F$, then we consider the exact complexity of $E$ to be completely determined. Short of doing this, we can establish upper bounds $E\leq_{B}F$ or $E<_{B}F$ for $E$, or lower bounds $F\leq_{B}E$ or $F<_{B}E$ for $E$, as much as we can, which provide partial information on the complexity of $E$.
We will use the following benchmark equivalence relations in this paper. Discussions about the ones without references can be found in [8].
  1. (1)
    $_{\mathbb{N}}$ is the equality relation on $\mathbb{N}$, which is a Polish space with the discrete topology. If $_{\mathbb{N}}\ \leq_{B}E$, then $E$ contains infinitely many equivalence classes.
  2. (2)
    $_{\mathbb{R}}$ is the equality relation on $2^{\mathbb{N}}$. If $_{\mathbb{R}}\ \leq_{B}E$, then $E$ contains continuum many equivalence classes.
  3. (3)
    $E_{0}$ is the eventual agreement relation on $2^{\mathbb{N}}$ defined as (x_i)_iE_0(y_i)_i⇔∃i ∀j≥i x_j=y_j.
  4. (4)
    $E_{\infty}$ is the universal countable Borel equivalence relation. It can be defined as the following equivalence relation on $2^{\mathbb{F}_{2}}$, where $\mathbb{F}_{2}$ is the free group with two generators: (x_g)_gE_∞(y_g)_g ⇔∃h∈F_2 ∀g∈F_2 x_g=y_gh.
  5. (5)
    $^{+}$ is the equivalence relation on $(2^{\mathbb{N}})^{\mathbb{N}}$ defined as (x_i)_i =^+ (y_i)_i⇔{x_i:i∈N}={y_i:i∈N}.
  6. (6)
    $^{++}$ is the equivalence relation on $((2^{\mathbb{N}})^{\mathbb{N}})^{\mathbb{N}}$ defined as (x_i)_i =^++ (y_i)_i⇔∀i ∃j x_i=^+ y_j and ∀j ∃i x_i=^+y_j.
  7. (7)
    1. a)
      AA is a hereditarily countable set in $\mathcal{P}^{3}(\mathbb{N})$;
    2. b)
      • for any $a,b\in A$ there is some $r$ such that $R(a,b,r)$ holds;
    The equivalence relation $\cong^{*}_{3,k}$ is defined as the isomorphism relation of countable structures coding pairs $(A,R)$ in $\mathcal{Q}^{*}_{k}(\mathbb{N})$. These equivalence relations were defined by Hjorth–Kechris–Louveau [14]. Our presentation above follows [2].
  8. (8)
    $E^{\infty}_{S_{\infty}}$ is the universal equivalence relation among all $S_{\infty}$-orbit equivalence relations, i.e., for any Borel action of $S_{\infty}$ on a standard Borel space $X$, the orbit equivalence relation $E^{X}_{S_{\infty}}=\{(x,y)\in X\colon\exists g\in S_{\infty}\;g\cdot x=y\}$ is Borel reducible to $E^{\infty}_{S_{\infty}}$. This equivalence relation is known to be universal among all isomorphism relations of countable structures. That is, given any Borel class of countable structures, the isomorphism relation is Borel reducible to $E^{\infty}_{S_{\infty}}$. Moreover, $E^{\infty}_{S_{\infty}}$ is also known to be Borel bireducible with many isomorphism relations of countable structures. For example, $E^{\infty}_{S_{\infty}}$ is Borel bireducible with the isomorphism relation for all countable linear orders.
The following Borel reducibility results are well known (see [8]):
$_{\mathbb{N}}\ <_{B}\ =_{\mathbb{R}}\ <_{B}E_{0}\,<_{B}E_{\infty}<_{B}\ =^{+}\,<_{B}\ =^{++}\ <_{B}E^{\infty}_{S_{\infty}}.$
It was also shown in Hjorth–Kechris–Louveau [14] that
$^{+}\,<_{B}\ \cong^{*}_{3,0}\ <_{B}\ \cong^{*}_{3,1}\ <_{B}\ =^{++}.$
The equivalence relations in (1) through (7) above are Borel equivalence relations, and $E^{\infty}_{S_{\infty}}$ is analytic and non-Borel.
Kechris–Nies–Tent [16] initiated the following research program.
Program.
  1. (1)
  2. (2)
Here $\mathcal{F}(S_{\infty})$ is the standard Borel space consisting of all closed subsets of $S_{\infty}$. The $\sigma$-algebra of Borel sets on $\mathcal{F}(S_{\infty})$ is generated by sets of the form
$\{F\in\mathcal{F}(S_{\infty})\colon F\cap U\neq\emptyset\},$
where $U\subseteq S_{\infty}$ is a basic open set, i.e., of the form
$N_{\bar{a},\bar{b}}=\{g\in S_{\infty}\colon g(\bar{a})=\bar{b}\}$
for some tuples $\bar{a},\bar{b}$ of natural numbers of the same length.
In this paper we will consider the class of closed subgroups of $S_{\infty}$ which are extremely amenable. We will show that it is a Borel subset of $\mathcal{F}(S_{\infty})$ (Theorem 6.1), thus this class fits well in the KNT research program.